# Low-Resolution Self-Attention for Semantic Segmentation

### v2

@@ Abstract
Semantic segmentation tasks naturally require high-resolution information for pixel-wise segmentation and global context information for class prediction. While existing vision transformers demonstrate promising performance, they often utilize high-resolution context modeling, resulting in a computational bottleneck. In this work, we challenge conventional wisdom and introduce the Low-Resolution Self-Attention (LRSA) mechanism to capture global context at a significantly reduced computational cost, i.e., FLOPs. Our approach involves computing self-attention in a fixed low-resolution space regardless of the input image’s resolution, with additional $3\times 3$ depth-wise convolutions to capture fine details in the high-resolution space. We demonstrate the effectiveness of our LRSA approach by building the LRFormer, a vision transformer with an encoder-decoder structure. Extensive experiments on the ADE20K, COCO-Stuff, and Cityscapes datasets demonstrate that LRFormer outperforms state-of-the-art models. The code is available at https://github.com/yuhuan-wu/LRFormer.
@@ 1 Introduction
Nevertheless, we observe that the computational overhead of self-attention remains a non-negligible bottleneck for existing vision transformers, as evidenced by Tab. XI. Consequently, we aim to delve deeper into the downsampling in the core component of the transformer, i.e., self-attention. Diverging from prior works that only downsample the key and value features [9, 23, 20], we propose to downsample all constituents—query, key, and value features. In this way, the output of self-attention would be in a low-resolution so that the mainstream of the transformer would contain low-resolution. Furthermore, we adopt a fixed downsampling size rather than a downsampling ratio to attain a very low computational complexity for self-attention. The proposed method is called Low-Resolution Self-Attention (LRSA).
@@
Figure 2: Experimental comparisons on ADE20K [27] dataset. Methods marked with $\dagger$” are the results pretrained on the ImageNet-22K dataset. Data are from Tab. III and Tab. XII.
@@ 2.1 Semantic Segmentation
Semantic segmentation is a fundamental task in computer vision. It is challenging due to the numerous variations like object sizes, textures, and lighting conditions in practical scenarios. FCN [30], the pioneering work in this area, proposed the adaptation of CNNs for semantic segmentation in an end-to-end manner. Since then, numerous studies have been built upon FCN [30], with major efforts focused on enriching multi-scale representations [4, 31, 5], enhancing boundary perception [32, 33, 34, 35], contextual representations [36, 21] and introducing visual attention[11, 3, 2, 14, 8]. These studies deeply explored the semantic head design upon FCN [30] and achieved great progress. Among these, many approaches [4, 1, 2, 3, 8, 5, 11, 12, 13, 14] are greatly benefited from the high-resolution features, performing prediction in the 1/8 of the input resolution to ensure high accuracy. Moreover, there exist some query-based frameworks [37, 38, 39, 39, 40, 41, 42] that changed the decoder style in semantic segmentation. For example, Panoptic SegFormer [37] introduced a unified mask prediction workflow using a query decoupling strategy to handle thing and stuff categories separately. MaskFormer [38] revolutionized mask decoders with transformer-based mask classification. Mask2Former [39] significantly enhanced MaskFormer via multi-scale masked attention. K-Net [40] unifies various segmentation tasks through dynamic kernels. K-means Transformer[41] integrated K-means clustering to optimize the mask generation for better instance
@@ 2.3 Vision Transformers
Transformers are initially proposed in natural language processing (NLP) [60]. Through multi-head self-attention (MHSA), transformers are capable of modeling global relationships. Thanks to this characteristic, transformers may also be powerful for computer vision tasks that require global information for a whole understanding of the visual scenarios. To bridge this gap, ViT [15] transformed an image to tokens via a 16$\times$16 pooling operation and adopts the transformer to process these tokens, achieving better performance than CNNs in image recognition. After that, vision transformers are developed rapidly by leveraging knowledge distillation [61], overlapping patch embedding [62] or convolutions [63, 64]. Recently, pyramid vision transformers [19, 23, 17, 10, 20, 26, 65, 66] are proven to be powerful for image recognition tasks like semantic segmentation. For example, PVT [19] and MViT [26] proposed to build a pyramid vision transformer pipeline via performing downsampling on key and value features. Notably, MViT [26] decreases the resolution of the query by half in the first block of each stage, without the need of patch embedding between each stage. Liu et al. [17] created a window-based vision transformer with shifted windows. Yuan et al. [10] presented HRFormer to learn high-resolution features for dense prediction using the vision transformer. Xia et al. [67] proposed DAT with deformable attention, conducting deformable sampling on key and value features. Wu et al. [20] introduced an efficient and multi-scale self-attention strategy via in-layer pyramid pooling. Liu et al. [68] proposed to compute the self-attention in a hierarchical manner.
Despite their reported effectiveness, it is still commonly believed that high-resolution features are crucial for self-attention to effectively capture contextual information in semantic segmentation. Window-based vision transformers [17, 24, 25, 18] calculate self-attention within each local windows to reduce the computational complexity so that they can keep the high-resolution of feature maps. Downsampling-based vision transformers [19, 23, 9, 26, 20, 68] keep the size of the query while partially conduct the downsampling on the key and value features with a fixed pooling ratio. Such strategy greatly reduces the complexity compared with vanilla attention so as to keep high-resolution features, making themselves computationally non-negligible especially for high-resolution inputs (Tab. XI). In contrast, we question the necessity of keeping high-resolution for capturing context information via self-attention. We study this question by proposing LRFormer with LRSA. The good performance on several public benchmarks suggest the superiority of our LRFormer for semantic segmentation.
@@ 3.1 Low-Resolution Self-Attention
where $d_{k}$ is the number of channels of $F_{in}$. We omit the multi-head operation for simplicity. The overall computational cost of vanilla self-attention is $O(N^{2}C+C^{2}N)$, where $N$ and $C$ are the number of tokens and the number of channels of $F_{in}\in\mathbb{R}^{N\times C}$, respectively. As the number of tokens of natural images are usually very large, the computational cost of vanilla self-attention is very high.
@@ 3.3 Decoder Head
As mentioned above, $F_{2},F_{3},F_{4}$ are resized to the same size as $F_{2}$ and then concatenated together. We apply a $1\times 1$ convolution on the concatenated feature to squeeze the number of channels. Then, a basic block (LRSA + FFN) is adopted to refine the squeezed feature. As we know, the feature from the top of the encoder, i.e., $F_{4}$, could be the most semantic meaningful. To avoid the loss of semantic information in the aggregation of high-level ($F_{4}$) and low-level ($F_{2},F_{3}$) features, we concatenate the refined feature with $F_{4}$ to enhance the semantics. After that, another basic block is connected for further feature refinement. Finally, we infer the segmentation prediction from the refined feature with a simple $1\times 1$ convolution. The experiments demonstrate that our simple decoder with LRSA can do better than previous state-of-the-art decoder heads for semantic segmentation, as shown in Tab. IX.
@@ 4.1 Experimental Setup
TABLE III: Comparisons with recent transformer-based methods on the ADE20K dataset [27]. The results of our method are marked as bold. “” indicates the result pretrained on ImageNet-22K.
@@ 4.2 Comparisons
ADE20K.  Results are shown in Tab. III. LRFormer is compared with several recent transformer-based methods in different complexity levels. Results of other methods are from their official repositories. We can observe that our LRFormer exhibits strong superiority over other methods. In terms of the mIoU, LRFormer-T/S/B/L are 4.5%/3.5%/2.6%/1.6% better than SegFormer-B1/B2/B4/B5 [17, 9]. LRFormer-T is 2.3% better than Swin-T-based Mask2Former [72] with near half FLOPs. With ImageNet-22K pretraining, LRFormer is 1.1% and 2.4% better than the strongest Swin-B-based MaskFormer [17, 38] and UperNet-based CSwin [81, 18] with fewer FLOPs. The visualized Fig. 2 of accuracy-FLOPs shows a more intuitive view of the comparisons.
@@
Cityscapes.  Tab. V presents the experimental comparisons between LRFormer and recent popular methods on the Cityscapes dataset. LRFormer outperforms SegFormer and HRFormer in all cases. We can observe that due to large input size, FLOPs of other methods are much higher than ours. For example, SegFormer-B2 costs 717G FLOPs while our LRFormer-S only spends 41% FLOPs with 0.9% improvement. More complexity analysis can refer to Tab. XI.
@@
ImageNet.  Since we pretrained our backbone encoder on ImageNet, we also evaluate our network on ImageNet classification only for reference. Results are shown in Tab. VI. We divide them to five groups. The four groups are divided by the FLOPs of approximate 2G, 4.5G, 9G, 16G, respectively. The fifth and sixth groups include the results are pretrained on ImageNet-22K dataset. The backbone encoder of our LRFormer outperformed recent state-of-the-art CNN-based methods such as ConvNeXt [56] and RepLKNet [57], and transformer-based methods like DAT [67] and P2T [20].
@@ 4.4 Ablation Study
Fixed pooled size.  We reported the results in Tab. VII for ADE20K semantic segmentation. For each basic block, the pooling operation will be omitted if the feature map size is smaller than the desired pooled size. Default fixed pooled size $m$ is $16^{2}$ for semantic segmentation. Results show larger pooled size ($m\geq 16^{2}$) achieves saturated performance. The default setting only introduces 5% training memory overhead and FLOPs compared with the pooled size of $8^{2}$ for semantic segmentation. Further decreasing the pooled size to $4^{2}$ will not introduce significant gain on improving efficiency. When increasing the pooled size to $32^{2},48^{2},64^{2}$, we obtain a minor improvement or even decreased performance on ADE20K semantic segmentation. We also observe that the FLOPs and training memory overhead are much more significant (26% $\sim$ 170%) when the pooled size is larger than $16^{2}$. We then conduct similar experiments on Cityscapes dataset, which has a much larger input size (1024$\times$1024). The mIoU results of LRFormer-L for pooled size of $16^{2}$ and $32^{2}$ are both 83.2%, showing that the default pooled size can also work well on larger input size. Considering the the performance, FLOPs and training memory, a low-resolution setting in LRFormer is much more proper.
@@
Locality capturing.  Our LRSA only computes the attention in low-resolution space. Introducing spatial locality, $3\times 3$ depth-wise convolution, to our network is beneficial for getting fine-grained semantic maps. In Tab. VIII, we analyzed the effect of the two depth-wise convolution before LRSA and in FFN. The number of GFLOPs for these three configurations is very similar so it is not reported. We can observe that the ADE20K performance of our LRFormer is improved by 0.5% and 1.4% and when adding the depth-wise convolution before LRSA in FFN, with 5% and 24% training memory overhead. If both depth-wise convolutions are removed, the mIoU performance will further drop 2.4% with 0.7GB less training memory usage, compared with the results only without the depth-wise convolution in FFN. This indicates that locality capturing plays a significant role in LRFormer. Therefore, we add both of them in our LRFormer.
@@
TABLE VIII: Ablation study on the spatial locality capturing. “Memory” is the training memory in ImageNet pretraining.
@@
TABLE IX: Comparisons of our simple decoder and other popular decoder heads.
@@
Comparions of different decoder heads.  Our decoder head aims to predict the semantic maps from multi-level feature maps effectively and efficiently with LRSA. The validate the LRSA of our decoder head, we compare it with several popular decode heads. These popular decoder heads are designed for CNNs, whose output feature maps are usually 1/8 of the original image. However, the backbone encoder of our LRFormer can output features of the 1/32 of the original image. To make a fair comparison, we first upsample features of the last stages and concatenate them together. Then, we feed them to the popular decoder heads. Other processes keep unchanged in these popular decoder heads. Tab. IX summarized the results on ADE20K semantic segmentation. The backbone is pretrained for 300 epochs on ImageNet-1K. Compared with PPM [4], DA [2], and CC [8], our LRFormer achieves 1.1%, 0.6%, and 0.9% improvement, respectively, with only fewer than 50% FLOPs. Compared with OCR [21], our LRFormer obtains 1.5% performance gain, with 83% FLOPs. Therefore, our default setting is more efficient and effective than other popular decoder heads.
@@
TABLE X: Discussions on the dimensions of the decoder. When the dimension of the decoder is larger than 384, the performance will be saturated or even decreased.
@@
TABLE XI: Analysis of the memory usage and FLOPs for different input size. “Att. FLOPs” indicates the summation of MHSA and upsampling operations. ”Memory” is the training memory for semantic segmentation.
@@
Dimensions of the decoder head.  To optimize the performance and computational cost of the decoder head, we employ a 1$\times$1 convolution to reduce the dimension of the concatenated multi-level features before feeding them into the decoder. We conducted experiments with various dimension settings and compared their results in Tab. X. The backbone is pretrained for 300 epochs on ImageNet-1K. The experiments show that a dimension setting of 512 achieves the best performance. However, setting the dimension to 384 results in only a 0.1% drop in mIoU performance, while saving 25% FLOPs. Therefore, we set the dimension of the decoder head to 384 in our LRFormer-S, reflecting the optimal trade-off between performance and computational cost.
Memory and FLOPs.  Our LRSA has a very low computational complexity of only $O(C^{2}+CN)$. We numerically analyze the efficiency of our LRFormer for different input sizes, as well as the comparisons with the representative method SegFormer [9]. The analyzed results on FLOPs, attention FLOPs, and training memory are shown in Tab. XI. For LRFormer, we additionally include the computational cost of upsampling operations. Our LRFormer-S costs much less memory and FLOPs than SegFormer-B2. Given input size of $1024\times 1024$, the number of FLOPs of MHSA operations in our LRFormer is dramatically lower than (0.9G vs. 54G) the self-attention in SegFormer. We can also observe that when the input size is increasingly larger, the superiority of LRFormer will be more substantial. This is because increasing input size will only slightly increase the FLOPs of upsampling operations in our MHSA.
TABLE XII: Comparisons with recent query-based frameworks on the ADE20K dataset [27]. The results of our method are marked as bold. Methods ended with “+” are the enhanced versions upgraded with Mask2Former’s decoder. “” indicates the result pretrained on ImageNet-22K and with a larger image size 640$\times$640.

### v3

@@ Abstract
Semantic segmentation tasks naturally require high-resolution information for pixel-wise segmentation and global context information for class prediction. While existing vision transformers demonstrate promising performance, they often utilize high-resolution context modeling, resulting in a computational bottleneck. In this work, we challenge conventional wisdom and introduce the Low-Resolution Self-Attention (LRSA) mechanism to capture global context at a significantly reduced computational cost, i.e., FLOPs. Our approach involves computing self-attention in a fixed low-resolution space, regardless of the input image’s resolution, with additional $3\times 3$ depth-wise convolutions to capture fine details in the high-resolution space. We demonstrate the effectiveness of our LRSA approach by building the LRFormer, a vision transformer with an encoder-decoder structure. Extensive experiments on the ADE20K, COCO-Stuff, and Cityscapes datasets demonstrate that LRFormer outperforms state-of-the-art models. Code is available at https://github.com/yuhuan-wu/LRFormer.
@@ 1 Introduction
Nevertheless, we observe that the computational overhead of self-attention remains a non-negligible bottleneck for existing vision transformers, as evidenced by Tab. XII. Consequently, we aim to delve deeper into the downsampling in the core component of the transformer, i.e., self-attention. Diverging from prior works that only downsample the key and value features [9, 23, 20], we propose to downsample all constituents—query, key, and value features. In this way, the output of self-attention would be in a low-resolution so that the mainstream of the transformer would contain low-resolution. Furthermore, we adopt a fixed downsampling size rather than a downsampling ratio to attain a very low computational complexity for self-attention. The proposed method is called Low-Resolution Self-Attention (LRSA).
@@
Figure 2: Experimental comparisons on ADE20K [27] dataset. Methods marked with $\dagger$” are the results pretrained on the ImageNet-22K dataset. Data are from Tab. III and Tab. XIII.
@@ 2.1 Semantic Segmentation
Semantic segmentation is a fundamental task in computer vision. It is challenging due to the numerous variations like object sizes, textures, and lighting conditions in practical scenarios. FCN [30], the pioneering work in this area, proposed the adaptation of CNNs for semantic segmentation in an end-to-end manner. Since then, numerous studies have been built upon FCN [30], with major efforts focused on enriching multi-scale representations [4, 31, 5], enhancing boundary perception [32, 33, 34, 35], contextual representations [36, 21] and introducing visual attention[11, 3, 2, 14, 8]. These studies deeply explored the semantic head design upon FCN [30] and achieved great progress. Among these, many approaches [4, 1, 2, 3, 8, 5, 11, 12, 13, 14, 37, 38, 39] are greatly benefited from the high-resolution features, performing prediction in the 1/8 of the input resolution to ensure high accuracy. Moreover, there exist some query-based frameworks [40, 41, 42, 42, 43, 44, 45] that changed the decoder style in semantic segmentation. For example, Panoptic SegFormer [40] introduced a unified mask prediction workflow using a query decoupling strategy to handle thing and stuff categories separately. MaskFormer [41] revolutionized mask decoders with transformer-based mask classification. Mask2Former [42] significantly enhanced MaskFormer via multi-scale masked attention. K-Net [43] unifies various segmentation tasks through dynamic kernels. K-means Transformer[44] integrated K-means clustering to optimize the mask generation for better instance
@@ 2.3 Vision Transformers
Transformers are initially proposed in natural language processing (NLP) [63]. Through multi-head self-attention (MHSA), transformers are capable of modeling global relationships. Thanks to this characteristic, transformers may also be powerful for computer vision tasks that require global information for a whole understanding of the visual scenarios. To bridge this gap, ViT [15] transformed an image to tokens via a 16$\times$16 pooling operation and adopts the transformer to process these tokens, achieving better performance than CNNs in image recognition. After that, vision transformers are developed rapidly by leveraging knowledge distillation [64], overlapping patch embedding [65] or convolutions [66, 67]. Recently, pyramid vision transformers [19, 23, 17, 10, 20, 26, 68, 69] are proven to be powerful for image recognition tasks like semantic segmentation. For example, PVT [19] and MViT [26] proposed to build a pyramid vision transformer pipeline via performing downsampling on key and value features. Notably, MViT [26] decreases the resolution of the query by half in the first block of each stage, without the need of patch embedding between each stage. Liu et al. [17] created a window-based vision transformer with shifted windows. Yuan et al. [10] presented HRFormer to learn high-resolution features for dense prediction using the vision transformer. Xia et al. [70] proposed DAT with deformable attention, conducting deformable sampling on key and value features. Wu et al. [20] introduced an efficient and multi-scale self-attention strategy via in-layer pyramid pooling. Liu et al. [71] proposed to compute the self-attention in a hierarchical manner. More approaches can refer to the survey [72].
Despite their reported effectiveness, it is still commonly believed that high-resolution features are crucial for self-attention to effectively capture contextual information in semantic segmentation. Window-based vision transformers [17, 24, 25, 18] calculate self-attention within each local windows to reduce the computational complexity so that they can keep the high-resolution of feature maps. Downsampling-based vision transformers [19, 23, 9, 26, 20, 71] keep the size of the query while partially conduct the downsampling on the key and value features with a fixed pooling ratio. Such strategy greatly reduces the complexity compared with vanilla attention so as to keep high-resolution features, making themselves computationally non-negligible especially for high-resolution inputs (Tab. XII). In contrast, we question the necessity of keeping high-resolution for capturing context information via self-attention. We study this question by proposing LRFormer with LRSA. The good performance on several public benchmarks suggest the superiority of our LRFormer for semantic segmentation.
@@ 3.1 Low-Resolution Self-Attention
where $d_{k}$ is the number of channels of $F_{in}$. We omit the multi-head operation for simplicity. The overall computational cost of vanilla self-attention is $O(N^{2}C+NC^{2})$, where $N$ and $C$ are the number of tokens and the number of channels of $F_{in}\in\mathbb{R}^{N\times C}$, respectively. As the number of tokens of natural images are usually very large, the computational cost of vanilla self-attention is very high.
@@ 3.3 Decoder Head
As mentioned above, $F_{2},F_{3},F_{4}$ are resized to the same size as $F_{2}$ and then concatenated together. We apply a $1\times 1$ convolution on the concatenated feature to squeeze the number of channels. Then, a basic block (LRSA + FFN) is adopted to refine the squeezed feature. As we know, the feature from the top of the encoder, i.e., $F_{4}$, could be the most semantic meaningful. To avoid the loss of semantic information in the aggregation of high-level ($F_{4}$) and low-level ($F_{2},F_{3}$) features, we concatenate the refined feature with $F_{4}$ to enhance the semantics. After that, another basic block is connected for further feature refinement. Finally, we infer the segmentation prediction from the refined feature with a simple $1\times 1$ convolution. The experiments demonstrate that our simple decoder with LRSA can do better than previous state-of-the-art decoder heads for semantic segmentation, as shown in Tab. X.
@@ 4.1 Experimental Setup
TABLE III: Comparisons with recent methods on the ADE20K dataset [27]. The results of our method are marked as bold. “” indicates the result pretrained on ImageNet-22K.
@@ 4.2 Comparisons
ADE20K.  Results are shown in Tab. III. LRFormer is compared with several recent transformer-based and Mamba-based methods in different complexity levels. Results of other methods are from their official repositories. We can observe that our LRFormer exhibits strong superiority over other methods. In terms of the mIoU, LRFormer-T/S/B/L are 4.5%/3.5%/2.6%/1.6% better than SegFormer-B1/B2/B4/B5 [17, 9]. LRFormer-T is 2.3% better than Swin-T-based Mask2Former [76] with near half FLOPs. With ImageNet-22K pretraining, LRFormer is 1.1% and 2.4% better than the strongest Swin-B-based MaskFormer [17, 41] and UperNet-based CSwin [86, 18] with fewer FLOPs. Compared with the representative Mamba-based Vim [79] with linear complexity, our LRFormer is significantly better. The visualized Fig. 2 of accuracy-FLOPs shows a more intuitive view of the comparisons.
@@
Cityscapes.  Tab. V presents the experimental comparisons between LRFormer and recent popular methods on the Cityscapes dataset. LRFormer outperforms SegFormer and HRFormer in all cases. We can observe that due to large input size, FLOPs of other methods are much higher than ours. For example, SegFormer-B2 costs 717G FLOPs while our LRFormer-S only spends 41% FLOPs with 0.9% improvement. More complexity analysis can refer to Tab. XII.
@@
ImageNet.  Since we pretrained our backbone encoder on ImageNet, we also evaluate our network on ImageNet classification only for reference. Results are shown in Tab. VI. We divide them to five groups. The four groups are divided by the FLOPs of approximate 2G, 4.5G, 9G, 16G, respectively. The fifth and sixth groups include results pretrained on ImageNet-22K dataset. The backbone encoder of our LRFormer outperformed recent state-of-the-art CNN-based methods such as ConvNeXt [59] and RepLKNet [60], and transformer-based methods like DAT [70] and P2T [20].
@@ 4.4 Ablation Study
Fixed pooled size.  We reported the results in Tab. VII for ADE20K semantic segmentation. For each basic block, the pooling operation will be omitted if the feature map size is smaller than the desired pooled size. Default fixed pooled size $m$ is $16^{2}$ for semantic segmentation. Results show larger pooled size ($m\geq 16^{2}$) achieves saturated performance. The default setting only introduces 5% training memory overhead and FLOPs compared with the pooled size of $8^{2}$ for semantic segmentation. Further decreasing the pooled size to $4^{2}$ will not introduce significant gain on improving efficiency. When increasing the pooled size to $32^{2},48^{2},64^{2}$, we obtain a minor improvement or even decreased performance on ADE20K semantic segmentation. We also observe that the FLOPs and training memory overhead are much more significant (26% $\sim$ 170%) when the pooled size is larger than $16^{2}$. We then conduct similar experiments on Cityscapes dataset, which has a much larger input size (1024$\times$1024). The mIoU results of LRFormer-L for pooled size of $16^{2}$ and $32^{2}$ are both 83.2%, showing that the default pooled size can also work well on larger input size. While adjusting pooled size based on input resolution could preserve more detail for small objects, our fixed-size design works well across all object scales. The high-resolution DWConv branch and multi-level feature aggregation effectively maintain small object information despite the spatial reduction. Furthermore, ideal backbone design paradigms [6, 59, 17, 19] suggest basic blocks maintaining consistent settings (same pooled size in our network) across different stages. This architectural simplicity enhances implementation efficiency and reduces the effort of finding optimal parameters for each stage individually. Considering the performance, FLOPs, training memory, and ideal backbone design paradigm, we use fixed low-resolution pooled size as the default setting.
@@
Locality capturing.  Our LRSA only computes the attention in low-resolution space. Introducing spatial locality, $3\times 3$ depth-wise convolution, to our network is beneficial for getting fine-grained semantic maps. In Tab. IX, we analyzed the effect of the two depth-wise convolution before LRSA and in FFN. The number of GFLOPs for these three configurations is very similar so it is not reported. We can observe that the ADE20K performance of our LRFormer is improved by 0.5% and 1.4% and when adding the depth-wise convolution before LRSA in FFN, with 5% and 24% training memory overhead. If both depth-wise convolutions are removed, the mIoU performance will further drop 2.4% with 0.7GB less training memory usage, compared with the results only without the depth-wise convolution in FFN. This indicates that locality capturing plays a significant role in LRFormer. Therefore, we add both of them in our LRFormer.
TABLE VIII: Performance comparison between the 16×16 pooled size (original) and smaller pooled size (4$\times$4).
  Category Metric Default Smaller Relative Change
  Small mIoU 36.2% 33.7% -7.1%
mAcc 46.3% 42.3% -8.8%
Medium mIoU 48.1% 45.3% -5.7%
mAcc 59.7% 56.8% -4.9%
Large mIoU 57.2% 53.8% -6.0%
mAcc 68.3% 64.3% -5.9%
Performance on small objects.  To investigate how pooled size affects different-sized objects, we categorize ADE20K semantic classes into small, medium, and large categories. Discrete objects were classified based on their typical real-world dimensions, while amorphous regions (e.g., sky) were included in the large category due to their typically extensive spatial coverage in images. Results are shown in Tab. VIII. Using a smaller pooled size (4$\times$4) instead of our default setting degrades performance across all categories, with small objects experiencing the most severe impact. This confirms that downsampling features to a resolution that is too small (e.g., 4$\times$4) can lose important semantics particularly for small objects.
@@
TABLE IX: Ablation study on the spatial locality capturing. “Memory” is the training memory in ImageNet pretraining.
@@
TABLE X: Comparisons of our simple decoder and other popular decoder heads.
@@
Comparions of different decoder heads.  Our decoder head aims to predict the semantic maps from multi-level feature maps effectively and efficiently with LRSA. The validate the LRSA of our decoder head, we compare it with several popular decode heads. These popular decoder heads are designed for CNNs, whose output feature maps are usually 1/8 of the original image. However, the backbone encoder of our LRFormer can output features of the 1/32 of the original image. To make a fair comparison, we first upsample features of the last stages and concatenate them together. Then, we feed them to the popular decoder heads. Other processes keep unchanged in these popular decoder heads. Tab. X summarized the results on ADE20K semantic segmentation. The backbone is pretrained for 300 epochs on ImageNet-1K. Compared with PPM [4], DA [2], and CC [8], our LRFormer achieves 1.1%, 0.6%, and 0.9% improvement, respectively, with only fewer than 50% FLOPs. Compared with OCR [21], our LRFormer obtains 1.5% performance gain, with 83% FLOPs. Therefore, our default setting is more efficient and effective than other popular decoder heads.
@@
TABLE XI: Discussions on the dimensions of the decoder. When the dimension of the decoder is larger than 384, the performance will be saturated or even decreased.
@@
TABLE XII: Analysis of the memory usage and FLOPs for different input size. “Att. FLOPs” indicates the summation of MHSA and upsampling operations. ”Memory” is the training memory for semantic segmentation.
@@
Dimensions of the decoder head.  To optimize the performance and computational cost of the decoder head, we employ a 1$\times$1 convolution to reduce the dimension of the concatenated multi-level features before feeding them into the decoder. We conducted experiments with various dimension settings and compared their results in Tab. XI. The backbone is pretrained for 300 epochs on ImageNet-1K. The experiments show that a dimension setting of 512 achieves the best performance. However, setting the dimension to 384 results in only a 0.1% drop in mIoU performance, while saving 25% FLOPs. Therefore, we set the dimension of the decoder head to 384 in our LRFormer-S, reflecting the optimal trade-off between performance and computational cost.
Memory and FLOPs.  Our LRSA has a very low computational complexity of only $O(C^{2}+CN)$. We numerically analyze the efficiency of our LRFormer for different input sizes, as well as the comparisons with the representative method SegFormer [9]. The analyzed results on FLOPs, attention FLOPs, and training memory are shown in Tab. XII. For LRFormer, we additionally include the computational cost of upsampling operations. Our LRFormer-S costs much less memory and FLOPs than SegFormer-B2. Given input size of $1024\times 1024$, the number of FLOPs of MHSA operations in our LRFormer is dramatically lower than (0.9G vs. 54G) the self-attention in SegFormer. We can also observe that when the input size is increasingly larger, the superiority of LRFormer will be more substantial. This is because increasing input size will only slightly increase the FLOPs of upsampling operations in our MHSA.
TABLE XIII: Comparisons with recent query-based frameworks on the ADE20K dataset [27]. The results of our method are marked as bold. Methods ended with “+” are the enhanced versions upgraded with Mask2Former’s decoder. “” indicates the result pretrained on ImageNet-22K and with a larger image size 640$\times$640.